A Database for the National Fetal and Infant Mortality Review Project

James R. Carbin

University of Rochester Medical Center

Abstract

The National Fetal and Infant Mortality Review Project (NFIMR) is a nation-wide project funded by the Robert Wood Johnson Foundation. One goal is to facilitate a systematic review of fetal and infant deaths which includes producing a Case Summary in text format. The other is to develop a National database of fetal/infant death. The focus: to identify causative factors which can lower infant/fetal mortality.

One NFIMR tool is the NFIMR Database System. Currently installed at 41 sites (15 states) with three states participating on a statewide basis, the requirements for this relational database system are complex. Extensive maternal and infant data as well as information about siblings (in the case of multiple gestations) is collected about each death from ten different sources. Potentially 3,872 different data items may be collected for one client. Data storage requires 42 files (tables).

This menu-driven system includes: a data dictionary and a User’s Guide (both are available in hard copy as well as through a context sensitive on line help system); encryption of sensitive data; passwords; extensive validity checking; and help screens to assist the user. The Case Summary is generated by parsing statements in a generic ”script” file. Medical personnel are responsible for writing these script files.

Refreshments and cookies will be served